# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v1

@@ IV.1 Pure ESS
$\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}$ $=-\epsilon_{y_{1}}\lambda^{\prime}_{1}-\epsilon_{y_{2}}\lambda^{\prime}_{2}-\cdots-\epsilon_{y_{n-1}}\lambda^{\prime}_{n-1}.$ (29)
We can see that the first $m-1$ eigenvalues are coming with a negation as the coefficient of the $\epsilon_{x_{i}}$’s in Eq. (28) and last $n-1$ eigenvalues are coming with a negation as the coefficient of $\epsilon_{y_{j}}$’s in Eq. (29). This indicates that for both equations, the RHS being positive for all neighborhood, implies, that all coefficients are positive. Consequently, all eigenvalues are negative, signifying dynamical stability. Analogously, one can demonstrate the validity of this statement for other corner fixed points. This demonstrate that a pure ESS is linearly stable fixed point if Definition 1d is used.
We also find that if we add Eq. (28) and Eq. (29) we get,
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-\sum_{i=1}^{m-1}\epsilon_{x_{i}}\lambda_{i}-\sum_{i=1}^{n-1}\epsilon_{y_{i}}\lambda^{\prime}_{i}.$ (30)
This implies that for a pure ESS under Definition 2d, in the entire local neighbourhood all the eigenvalue must be negative. Thus, in the context of linear stability, Definitions 1d and 2d appear equivalent.
Even more importantly, in the light of Eq. (28) and Eq. (29), Definition 3d implies
either $\text{either }-\sum_{i=1}^{m-1}\epsilon_{x_{i}}\lambda_{i}>0~{}~{}\text{or}~{}~{}-\sum_{i=1}^{n-1}\epsilon_{y_{i}}\lambda^{\prime}_{i}>0,$ (31)
whose validity in the entire neighbourhood does not guarantee the stability of the corresponding fixed point because it allows for some positive eigenvalues. Therefore, one may find payoff matrices for which Definition 3d is satisfied, but the corresponding fixed point is not stable, thus contradicting Theorem 1. For example, if we consider the following payoff matrices:
${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}\frac{1}{3}&\frac{1}{3}\\ 0&\frac{2}{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},$ (32)
we can easily see that the corner fixed point $(1,1)$ of the replicator dynamics, satisfies Definition 3d: $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\tfrac{1}{12}\epsilon_{x}(1+9\epsilon_{x}-8\epsilon_{y})>0$. However, the set of eigenvalues $\{\lambda_{1},\lambda^{\prime}_{1}\}=\{-\tfrac{1}{12},\tfrac{7}{12}\}$ indicate that the fixed point is not stable as further depicted in Fig. 1(b). One must be aware of the subtlety that even with some eigenvalues zero and some negative, inequality condition (31) may be satisfied. In such a case, as linear stability analysis would be deemed a failure, maybe going to nonlinear terms would yield a pure 2ESS that is locally asymptotically stable. As an example, consider the following
${\sf A}=\begin{pmatrix}1&\frac{1}{3}\\ 3&\frac{1}{3}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}0&\frac{1}{3}\\ 1&\frac{1}{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}2&\frac{1}{3}\\ 1&\frac{1}{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}1&\frac{1}{3}\\ 2&\frac{1}{3}\end{pmatrix}.$ (33)
We can easily see the corner fixed point $(0,0)$ of the replicator dynamics is a 2ESS that is, it satisfies Definition 3d: $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\epsilon_{x}(\epsilon_{y}+2\epsilon_{x})>0$. To see that it only satisfies the weaker Definition 3d and not Definition 1d, we note, $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=\epsilon_{y}(\epsilon_{y}-\epsilon_{x})$ has presence of the negative sign. The eigenvalues are both zero in this case. We can see form Fig. 1(c) that $(0,0)$ is indeed an asymptotically stable fixed point.
With the above results in the backdrop, one can go beyond linear stability analysis. Consider the following function:
Evidently, realizing that the two sums in $V_{1}$ represent Kullback–Leibler (KL) divergence Kullback and Leibler (1951); Cover and Thomas (2005) of $\mathbf{\hat{x}}$ from $\mathbf{x}$ and KL-divergence of $\mathbf{\hat{y}}$ from $\mathbf{y}$, $V_{1}$ must be positive except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $V_{1}=0$. If additionally we could show that $\dot{V}_{1}<0$ $\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then $V_{1}$ qualifies as Lyapunov function Jordan and Smith (1999); Wiggins (2003) and the ESS must be locally asymptotic fixed point. To this end, on taking the time derivative, we get,
@@
where we have used Eq. (26). Of course, $\dot{V_{1}}<0$ $\forall({\mathbf{x}},{\mathbf{y}})\neq({\mathbf{\hat{x}}},{\mathbf{\hat{y}}})$ when either inequalities (23) or inequalities (24) are satisfied. Hence, any pure ESS under either Definition 1d or Definition 2d has $V_{1}$ as a Lyapunov function and the corresponding fixed point is locally asymptotically stable. These conclusions carry over to bimatrix games (${\sf A}={\sf D}={\sf 0}$) as well: The same Lyapunov function can be used to show that any pure ESS under Definition 1c or 2c is locally asymptotically stable.
IV.1.1 A Caveat: Definition 1 versus Definition 2
Note, however, that it is not necessary that an ESS under Definition 2d is always an ESS under Definition 1d. A general class of such ESS’s can be found as the cases of non-strict pure Nash equilibria in bimatrix games. Explicitly, consider
where $b_{1}>0$ and $c_{3}>0$. We can see that $(\hat{x},\hat{y})=(1,1)$ is a non-strict Nash equilibrium. To locate ESS, just like in Section III, we denote the probability vector corresponding to the non-strict Nash as $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ and the neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3})$ which is positive for infinitesimally small $\epsilon$ and $\delta$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta\epsilon c_{3}$ which is negative—we conclude that $(\hat{x},\hat{y})$ is a 2ESS. But it does not satisfy Definition 1c. To find whether it satisfies Definition 2c of ESS, we calculate: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3}-c_{3})$. It means $(\hat{x},\hat{y})$ satisfies Definition 2c. What about its stability?
To see the dynamical stability, we write the replicator equation for the payoff matrix ${\sf B}$ and ${\sf C}$ from Eq. (36):
$\dot{x}$ $=$ $x(1-x)\left[b_{2}-b_{3}+y(b_{1}-b_{2}+b_{3})\right],$ (37)
$\dot{y}$ $=$ $-y(1-y)(1-x)c_{3}.$ (38)
We can see that the fixed point $(1,1)$ is not asymptotically stable because $\dot{y}$ is always negative causing $y$-value to always decrease away from $1$. Consequently, phase points near $(1,1)$ moves away from it, making it an unstable fixed point. This demonstrates that, in the case of non-strict Nash equilibria, Definition 1c is a better indicator of stability than Definition 2c. This is because failing to satisfy Definition 1c renders the fixed point to be unstable while ESS satisfying Definition 2c wrongly predicts the corresponding fixed point to be stable. This conclusion is not in contradiction with the equivalence between the two definitions in the context of linear stability and Lyapunov function. This is because there the linear stability analysis fails as one eigenvalue is zero and the Lyapunov function argument is not applicable since the fixed point is not isolated—$x=1$ is a line of non-isolated fixed points.
IV.2 Mixed ESS
The case of mixed ESS is of relevance in asymmetric game as now it in principle could be attracting—which is not possible in bimatrix case—under the replicator dynamics on the phase space. For illustration, on choosing $m=n=2$, consider the internal fixed point,
$\left(\hat{x},\hat{y}\right)=\left(\frac{-a^{\prime}d+bc^{\prime}+bd^{\prime}-b^{\prime}d}{ad-bc},\frac{-ac^{\prime}-ad^{\prime}+a^{\prime}c+b^{\prime}c}{ad-bc}\right),$ (39)
if it exists in the sense that expressions are real numbers between zero and one. Here, $a\equiv a_{11}-a_{12}-a_{21}+a_{22}$$a^{\prime}\equiv a_{12}-a_{22}$$b\equiv b_{11}-b_{12}-b_{21}+b_{22}$$b^{\prime}\equiv b_{12}-b_{22}$$c\equiv c_{11}-c_{12}-c_{21}+c_{22}$$c^{\prime}\equiv c_{12}-c_{22}$$d\equiv d_{11}-d_{12}-d_{21}+d_{22}$ and $d^{\prime}=d_{12}-d_{22}$. On linearization, the eigenvalues, say, $\lambda_{-}$ and $\lambda_{+}$, of the Jacobian matrix comes out as
It is evident that now there is a possibility of focus or node types of fixed points and hence, under suitable conditions, if there is a mixed ESS, then its corresponding phase point may be a locally asymptotic fixed point.
Let us now consider the general case of mixed ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ when $m$ and $n$ are arbitrary. This should be fixed point of the corresponding replicator equation. By examining the same sum of KL-divergence, $V_{1}(\mathbf{x},\mathbf{y})$ from Eq. (34), we realize that a mixed ESS under either Definition 1d or Definition 2d has $V_{1}$ as a Lyapunov function, making the corresponding fixed point locally asymptotically stable. It is important to be reminded of that there can be an ESS under Definition 2d even when there is none under Definition 1d as the former is a weaker definition. Now what about the even weaker definition of ESS, viz., Definition 3d?
A theorem Cressman (1996) comes to our help:
@@ IV.2 Mixed ESS
qualifies as the Lyapunov function for a mixed 2ESS which, thus, is locally asymptotically stable fixed point. Again, we note that while ESS under Definitions 1d and 2d must be 2ESS, the converse need not be true. There is a subtlety to be noted: This existence of this Lyapunov function does not extend to bimatrix games (see Appendix A) formed by setting ${\sf A}={\sf 0}$ and ${\sf D}={\sf 0}$.
Finally, all the considerations up to now lead us to believe that, keeping in view the desirable dynamical aspects of evolutionary stability, Definition 1d is the most compatible and natural extension of the concept of ESS to the case of asymmetric and bimatrix games. While unlike this definition, 2ESS (equivalently, Nash–Pareto pair) may allow for mixed ESS in bimatrix games, the mixed ESS cannot be asymptotically stable and hence is not of much practical consequence in our view. As far as, the pure ESS is concerned in bimatrix games, 2ESS does not guarantee its asymptotically stability. Only relevance of 2ESS might be in asymmetric games, where it is surely locally asymptotically stable as interior fixed point, because 2ESS may allow for mixed ESS that would be otherwise be absent under Definition 1d and 2d.

### v2

@@ IV.1.1 Strict NE as ESS
$\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}$ $=-\delta_{y_{1}}\lambda^{\prime}_{1}-\delta_{y_{2}}\lambda^{\prime}_{2}-\cdots-\delta_{y_{n-1}}\lambda^{\prime}_{n-1}.$ (37)
We can see that the first $m-1$ eigenvalues, as the coefficient of the $\delta_{x_{i}}$’s, come with negative signs in Eq. (36). Similarly, last $n-1$ eigenvalues, as the coefficient of $\delta_{y_{j}}$’s, come with negative signs in Eq. (37). If the fixed point has to be ESS then for both equations, the RHS’s have to be positive in the entire neighbourhood. This requirement equivalently demands that all the eigenvalues be negative because all the $\delta$’s are independent of each other and all are not allowed to be simultaneously zero. This demand is automatically satisfied if the fixed point is strict NE.
We furthermore find that if we add Eq. (36) and Eq. (37) we get,
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}-\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}.$ (38)
The LHS has to be positive for the fixed point to be ESS under sum-definition. It is easy to see that exactly same argument as given above for the ESS under and-definition applies to this case. Hence, in the context of linear stability, and- and sum-definitions appear equivalent.
Lastly, in the light of Eq. (36) and Eq. (37), or-definition implies
either $\text{either }-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}>0~{}~{}\text{or}~{}~{}-\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}>0,$ (39)
for the fixed point to be ESS. The validity of condition (39) in the entire neighbourhood requires the negativity of all the eigenvalues, or equivalently, requires the fixed point to be strict NE. It might be tempting to assume that not all eigenvalues need to be negative for condition (39) to hold, given the presence of conjunction or. However, it is important to recall that in Definition 3d, mutants in either or both subpopulations are allowed, as discussed in Section II.2. Therefore it is allowed to have either all $\delta_{x}$’s or all $\delta_{y}$’s to be zero rendering either the former or the latter inequality in condition (39) invalid. As a result, the inequalities in (39) imply that all eigenvalues must be negative. In conclusion, strict NE is 2ESS as well.
IV.1.2 Non-strict NE as ESS
The equivalence between ESS’s under and-, sum- and or-definitions goes away if the ESS is non-strict NE. Following example highlights this.
Example 4 (see Fig. 1d): Non-strict pure NE as ESS under or-definition need not be ESS under and- and sum-definitions.
Consider
${\sf A}=\begin{pmatrix}a_{1}&a_{2}\\ a_{1}&a_{3}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}d_{1}&d_{2}\\ d_{1}&d_{3}\end{pmatrix},$ (40)
where $a_{2}>a_{3}$, $b_{2}>b_{3}$, $c_{2}<c_{3}$ and $d_{2}>d_{3}$. We can see that $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is a non-strict NE whose neighbourhood state is given as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. To check whether it is an ESS we calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\epsilon\delta(b_{2}-b_{3})+\epsilon^{2}(a_{2}-a_{3})$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=\epsilon\delta(c_{2}-c_{3})+\delta^{2}(d_{2}-d_{3})$. It is an ESS under or-definition since at least one of them is positive for all allowed values of $\epsilon$ and $\delta$. However, it is not an ESS under and-definition since $c_{2}-c_{3}<0$; furthermore, if $|c_{2}-c_{3}|>|b_{2}-b_{3}|$, then it may not satisfy an ESS condition under sum-definition. The fixed point remains asymptotically stable as may be checked numerically, thus, validating Theorem $1^{\prime}$.
Before going to the case of mixed ESS, we would like to bring it to the fore that for Definitions 1d and 2d, there are further arguments based on Lyapunov function for them to obey Theorem 1 in case of pure ESS corresponds to an isolated fixed point. One can show using the same Lyapunov function given in Eq. (20), that satisfying Definition 1d or 2d imply asymptotic stability for the pure ESS (corner fixed point of the simplotope). However, the general proof that Definition 3d imply asymptotic stability for pure ESS is not so straightforward: A proof using center manifold theory is available in the literature Cressman (1996).
IV.2 Mixed ESS
The introduction of intra-specific interactions leads to circumventing certain restrictions present in bimatrix games. Specifically now and- and sum-definitions can lead to mixed ESS and mixed ESS under or-definition can be locally asymptotically stable. While the former is illustrated in Example 5 below, the latter is easily seen on choosing $m=n=2$, in which case the form of internal fixed point, it it exists, is
$\left(\hat{x},\hat{y}\right)=\left(\frac{-a^{\prime}d+bc^{\prime}+bd^{\prime}-b^{\prime}d}{ad-bc},\frac{-ac^{\prime}-ad^{\prime}+a^{\prime}c+b^{\prime}c}{ad-bc}\right).$ (41)
Here, $a\equiv a_{11}-a_{12}-a_{21}+a_{22}$$a^{\prime}\equiv a_{12}-a_{22}$$b\equiv b_{11}-b_{12}-b_{21}+b_{22}$$b^{\prime}\equiv b_{12}-b_{22}$$c\equiv c_{11}-c_{12}-c_{21}+c_{22}$$c^{\prime}\equiv c_{12}-c_{22}$$d\equiv d_{11}-d_{12}-d_{21}+d_{22}$ and $d^{\prime}=d_{12}-d_{22}$. On linearization, the eigenvalues, say, $\lambda_{-}$ and $\lambda_{+}$, of the Jacobian matrix comes out as
It is evident that now there is a possibility of focus or node types of fixed points under suitable conditions. For clear visualization, consider the example of asymmetric game given in (44), where change in the stability of the fixed point—due to addition of intra-specific interaction—is demonstrated through phase portraits as plotted in Fig. 1(c) and Fig. 1(f). Furthermore, Example 6 (see below) makes an important point clear that 2ESS may help find new evolutionarily important asymptotically stable states which may escape and- and sum-definitions.
Example 5 (see Fig. 1e): Mixed ESS is possible under and- and sum-definitions in asymmetric games.
Suppose that
${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (43)
One can easily check that the internal fixed point (mixed NE) in this case is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$. Therefore the neighbourhood vector has the form—$(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}{2}/{9}+\delta_{x},{7}/{9}-\delta_{x}\end{pmatrix},\begin{pmatrix}{2}/{9}+\delta_{y},{7}/{9}-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small and can be positive or negative but they can’t be zero simultaneously. We calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}={3\delta_{x}^{2}}/{4}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}={3\delta_{y}^{2}}/{4}$. Both terms are positive for all infinitesimal values of $\delta_{x}$ and $\delta_{y}$, therefore it satisfies and-definition. For sum-definition one just need to add both the terms and get, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}={3(\delta_{x}^{2}+\delta_{y}^{2})}/{4}$, which means sum-definition is also satisfied. Fig. 1(e) shows that the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$ is asymptotically stable, confirming Theorem 1 for and- and sum-definitions.
Example 6 (see Fig. 1f): Mixed 2ESS need not be ESS under and- and sum-definitions.
Consider the payoff matrices,
${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}1&4\\ 2&2\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}6&4\\ 3&8\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (44)
Now, let see whether the internal fixed point is an ESS. One can calculate the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}66/115,49/115\end{pmatrix},\begin{pmatrix}238/345,107/345\end{pmatrix}\right)$, the neighbourhood vector therefore is, $(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}\frac{66}{115}+\delta_{x},\frac{49}{115}-\delta_{x}\end{pmatrix},\begin{pmatrix}\frac{238}{345}+\delta_{y},\frac{238}{345}-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small. We calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=3\delta_{x}\delta_{y}+\frac{3\delta_{x}^{2}}{4}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-7\delta_{x}\delta_{y}+\frac{3\delta_{y}^{2}}{4}$. It is not very hard observe that at least one of the terms in the RHS of the two expressions is positive for all possible infinitesimal values of $\delta_{x}$ and $\delta_{y}$, which means the internal fixed point is 2ESS—an ESS under or-definition. Since both the terms cannot remain simultaneously positive in the entire neighbourhood, it is definitely not an ESS under and-definition. To check for sum-definition we add the two terms and get, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-4\delta_{x}\delta_{y}+\frac{3(\delta_{x}^{2}+\delta_{y}^{2})}{4}$, which is not guaranteed to be always positive in the neighbourhood; hence it is not an ESS under sum-definition as well. Fig. 1(f) shows that the fixed point, a 2ESS, is asymptotically stable, confirming Theorem 1 for or-definition.
The stability of the mixed ESS can be understood through Lyapunov function. Let us consider the general case of mixed ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ when $m$ and $n$ are arbitrary by considering the following function given in Eq. (20). We calculate the time derivative using Eq. (34) and get,
@@ IV.2 Mixed ESS
We can see that $\dot{V_{1}}<0$ $\forall({\mathbf{x}},{\mathbf{y}})$ in some neighbourhood of $({\mathbf{\hat{x}}},{\mathbf{\hat{y}}})$ when either inequalities (31) or inequalities (32) are satisfied. Hence, any mixed ESS under either and-definition or sum-definition has $V_{1}$ as a Lyapunov function and the corresponding fixed point is locally asymptotically stable. Note that, these conclusions does not carry over to bimatrix games (${\sf A}={\sf D}={\sf 0}$) for mixed ESS since we know only pure ESS is possible there. As far as the or-definition is concerned, a theorem Cressman (1996) comes to our help:
@@
qualifies as the Lyapunov function for a mixed 2ESS which, thus, is locally asymptotically stable fixed point. There is a subtlety to be noted: This existence of this Lyapunov function does not extend to bimatrix games (see Appendix A) formed by setting ${\sf A}={\sf 0}$ and ${\sf D}={\sf 0}$.